Due to erratic sales of its sole product—a high-capacity battery for laptop computers—PEM, Inc., has been experiencing financial difficulty for some time. The company’s contribution format income statement for the most recent month is given below:

  

Sales (13,500 units × $30 per unit) $ 405,000
Variable expenses 202,500
Contribution margin 202,500
Fixed expenses 225,000
Net operating loss $ (22,500 )

Required:

1. Compute the company’s CM ratio and its break-even point in unit sales and dollar sales.

2. The president believes that a $6,800 increase in the monthly advertising budget, combined with an intensified effort by the sales staff, will result in an $85,000 increase in monthly sales. If the president is right, what will be the increase (decrease) in the company’s monthly net operating income?

3. Refer to the original data. The sales manager is convinced that a 10% reduction in the selling price, combined with an increase of $39,000 in the monthly advertising budget, will double unit sales. If the sales manager is right, what will be the revised net operating income (loss)?

4. Refer to the original data. The Marketing Department thinks that a fancy new package for the laptop computer battery would grow sales. The new package would increase packaging costs by 0.60 cents per unit. Assuming no other changes, how many units would have to be sold each month to attain a target profit of $4,400?

5. Refer to the original data. By automating, the company could reduce variable expenses by $3 per unit. However, fixed expenses would increase by $57,000 each month.

a. Compute the new CM ratio and the new break-even point in unit sales and dollar sales.

b. Assume that the company expects to sell 20,900 units next month. Prepare two contribution format income statements, one assuming that operations are not automated and one assuming that they are. (Show data on a per unit and percentage basis, as well as in total, for each alternative.)

c. Would you recommend that the company automate its operations (Assuming that the company expects to sell 20,900)?
Answer #1
contribution margin per unit= 202500/13500
15
1) CM ratio = contribution/sales
202500/405000
50.00%
BEP(units) = total fixed cost/contribution margin per unit
225000/15
15000
BEP(dollars) = 15000*30
450000
CM ratio 50%
Break even point in units 15000
Break even point in dollars 450000
2) increase in contribution (85000*50%) 42500
less : increase in advertising budget 6,800
increase in net income 35,700
increases by 35,700
3) units = 13500*2 = 27000 units ; selling price = 30*90%=$27
Contribution Income statement
Sales (27000*27) 729000
Variable expense (27000*15) 405000
Contribution margin 324000
Fixed expenses (225000+39000) 264,000
Net income 60,000
4) New contribution margin = 15-.60
14.4
BEP(units) = (total fixed cost+target profit)/contribution per unit
(225000+4400)/14.4
15930.56
Sales units 15,931
5)
CM ratio = contribution/sales
18/30
60.00%
BEP(units) = total fixed cost/contribution margin per unit
(225000+57000)/18
15667
BEP(dollars) = 282000/60%
470000
CM ratio 60%
Break even point in units 15667
Break even point in dollars 470000
20900
b) Not Automated Automated
total per unit % total per unit %
Sales 627000 30 100% 627000 30 100%
Variable expenses 313500 15 60% 250800 12 50%
Contribution margin 313500 15 40% 376200 18 50%
Fixed expenses 225,000 282,000
Net operating income 88,500 94,200
c) yes